BRIDGING THE GAP BETWEEN VAN DER WAALS AND HYDROGEN BONDING: A NEAR-IR STUDY OF THE INTERMOLECULAR MODES IN HCl DIMER

M. FARNIK, \underline{S. DAVIS}, and D. J. NESBITT, JILA, National Institute of Standards and Technology and Department of Chemistry and Biochemistry, University of Colorado, Colorado 80309-0440.

The near IR combination band spectra of supersonically cooled (HCL)_{2} in the 2900 cm^{−1} to 3070 cm^{−1} region have been recorded with a high resolution slit jet spectrometer. Seven vibration-rotation-tunneling (VRT) bands are observed, representing 3 (out of 4 total) intermolecular modes (van der Waals stretch v_{4}, geared bend v_{5}, and out-of-plane torsion v_{6}) built as combination bands on either the v_{1} (free) or v_{2} (bound) HCl stretches. Each of the seven combination bands are observed for both the H^{35}Cl-H^{35}Cl and H^{35}Cl-H^{37}Cl isotopomers. Analysis of the rotationally resolved spectra provide spectroscopic constants, intermolecular frequencies, predissociation rates, and isotopic shifts as a function of both intraand inter-molecular excitation. The intermolecular frequencies are compared with full 6-D quantum calculations of a trial potential surface and discrepencies are discussed. Additionally, comparisons with similar studies on (HF)_{2} and (DF)_{2} are presented. Both tunneling rates and combination band intensities are compared between the chlorine and fluorine build dimers, illuminating dramatic differences in these seemingly similar hydrogen bonded prototypes.